OpenStudy (anonymous):

A factory can produce two products, x and y, with a profit approximated by P = 14x + 22y – 900. The production of y can exceed x by no more than 100 units. Moreover, production levels are limited by the formula x + 2y ≤ 1,400. What production levels yield maximum profit?

OpenStudy (anonymous):

@zaynahf

OpenStudy (anonymous):

Ok I think I finally solved this one, and to be honest I despise these types of questions :) but here goes anyways. With you wanting the maximum profit defined by P=14x+22y-900, which parameter has the maximum effect on profit? Since y is multiplied by 22 you really want to maximize how much y shows up to make much moolah. Now we move onto x+2y<=1400, again, since we want the max we won't use less that 1400 total.. that would be silly, since what are we looking for? moolah. Now finding a relationship between x and y was my next step, because I want to find just how much y I can have without going over 1400. The best would just be to use y for the formula, but x is tied to it with a margin of 100, but again we want to minimize x's involvement so we can deduce the following relationship: x=y-100. Back to x+2y<=1400. Now we can plug in our relationship to just have a single variable equation and solve. y=500 is what we find. Thus, x=400. I'm quite confident this is correct.

OpenStudy (amoodarya):

look at image below

OpenStudy (anonymous):

That looks really nice, but it far exceeds the limit of 1400<=x+2y

OpenStudy (amoodarya):

look now

OpenStudy (anonymous):

that still exceeds the limit :/

OpenStudy (amoodarya):

plz write all the limits in math